# In-context Time Series Predictor

### v1

@@ 3 Experiments
We conduct comprehensive experiments under full-data, few-shot, and zero-shot settings using widely-used TSF datasets (details in §A.2), including ETTs [16], Traffic, Electricity (ECL), and Weather. We use $K=3$ $\mathtt{TF}$ layers with $d=128$ and 8 heads. We set $L_{I}=1440$, $L_{b}=512$, and $L_{P}\in\{96,192,336,720\}$, performing 4 experiments for each dataset. We use sampling step $m=8$ and the token retrieval method with $q=10\%$, $r=30$ in main experiments. See §A.3.2 for more hyper-parameter and implementation details. We use the same experimental environment as [25, 26]. Extensive efforts are made to ensure fair comparisons across all experiments (detailed in §A.3.3).
Baselines We compare our method with previous SOTA models in the categories of LLM for TSF (Time-LLM [26], GPT4TS [25], LLMTime [24]), Temporal-wise Transformers (PatchTST [18], FEDformer [28], Autoformer [17], Informer [16]), Series-wise Transformer (iTransformer [21]), CNN (TimesNet [29]), and simple methods (DLinear [19], Last-value Repeat). Baseline results are sourced from [25, 26] when applicable and rerun for the missing experiments. We report the average test set MSE for each dataset in the main paper and provide the full raw results in §A.1. The average rank and the number of times each model is ranked as the best are also summarized in the results.
@@
Analysis of TSF Transformer Characteristics From the ICL perspective, the Temporal-wise Transformer is better for learning inter-series relationships, while the Series-wise Transformer better at intra-series relationships. Here, we use a synthesized "Multi" dataset from [20] with strong inter-series dependencies, generated by shifting an random walk process. Fig. 4 (a) and (b) show that the Temporal-wise Transformer performs well on this dataset, while the Series-wise Transformer fails to learn the shifting dependencies. ICTSP, with its context tokens featuring a shifting structure, excels in learning these time-interleaved dependencies. Conversely, the Temporal-wise Transformer struggles with real-world datasets with weak dependencies comparing to Series-wise Transformer, such as the ETTm2 datasets in Fig. 4 (c) and (d). ICTSP handles these datasets well, demonstrating significantly better performance. Note that we build these baselines in this analysis with the same structure shown in Fig. 2, using $L_{I}=512$, $L_{P}=192$, $K=3$, $d=128$ for the 3 models; and $L_{b}=256$ for ICTSP. See §A.4.2 for more discussion about the inter-series learning abilities of the TSF Transformers.
@@ 4 Related Works
Extensive research on TSF has been conducted in both traditional statistics [30, 31, 32] and deep learning [33, 34, 35]. Recent studies have seen a surge in Transformer-based TSF models [36, 15, 16, 17, 28, 37]. However, a study [19] indicated that these complex models might not outperform linear predictors on real-world datasets, leading to subsequent research [18, 21, 20] proposing possible solutions. Additionally, due to the significant achievements of LLMs in various tasks [1], several LLM-based TSF models [24, 25, 26] have been introduced. These models also use Transformer architectures but rely on pre-trained parameters and fine-tuning to achieve superior few-shot and zero-shot capabilities [23]. While studies [2, 3, 4, 5, 6] have highlighted that these few-shot abilities in LLMs stem from ICL, there is a lack of research on how to efficiently implement ICL in TSF.
@@ A.2 Datasets
ETT Datasets [^1] https://github.com/zhouhaoyi/ETDataset [16]: This dataset includes load and oil temperature data from electricity transformers, recorded at 15-minute intervals from July 2016 to July 2018. It comprises four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, representing two transformers (identified as 1 and 2) and two time resolutions (15 minutes and 1 hour). Each subset contains seven features related to oil and load of the transformers.
@@
Multi Dataset [20] is generated from a master random walk series. The first series is the master series, while the second to fifth series are created by shifting the master series backward by 96, 192, 336, and 720 steps, respectively. The last three series are combinations of these first five series.
@@ A.3.1 More Details of Model Implementation
The remaining tokens are merged into $r$ tokens by grouped weighted averaging. Let $k=\lfloor qN\rfloor$ be the number of selected tokens, and let $p=N-k$ be the number of remaining tokens. The $r$ merged tokens are obtained as follows:
@@ A.3.2 Hyperparameter Settings and Training Details
Model Training Setup The ICTSP model is trained using the Adam optimizer and MSE loss in Pytorch, with a learning rate of 0.0005 each dataset. We test the model every 200 training steps with a early-stopping patience being $30\times 200$ steps. The first 1000 steps are for learning rate warm-up, followed by a linear decay of learning rate. We set the random seed as 2024. Our models are trained on single Nvidia RTX 4090 GPU with a batch size equals to 32 for most of the datasets. We decrease the batch size to 16 and 8 for the larger dataset ECL and Traffic, respectively. In the full-data experiment setting, we split each dataset with 70% training set, 10% validation, set and 20% test set. We fit a standardization scaler on the training set and apply it to the whole dataset. This setting is applied in consistent with previous studies like [17, 29, 18, 26].
@@
Randomness of Training We test the randomness of training the ICTSP by alternating the random seed among {2022, 2023, 2024, 2025, 2026} and conducting multiple training runs on the same dataset settings. The error bars of the results with the five random seeds on the Weather and ECL datasets are shown in Figure 8, alongside the results of the two best previous baselines, Time-LLM [26] and PatchTST [18]. The results show that ICTSP can stably surpass previous methods, even with the worst results among the five random seeds, indicating the robustness of ICTSP in performance improvement.
@@ A.3.3 Discussion of Fair Comparison
Our experimental environment aligns with the codes used by previous studies, including [17, 19, 29, 18, 25, 26]. Our few-shot and zero-shot learning setups follow [25, 26].
In the full-data experiment, for the input length $L_{I}$ setting, previous studies like [18, 25] suggest that $L_{I}$ should be considered a hyper-parameter that needs tuning for each dataset. Therefore, we chose to directly source their results where applicable, to avoid underestimating their performance when rerunning with a fixed $L_{I}$.
In the few-shot experiment, careful consideration of $L_{I}$ is very crucial. Since [25, 26] use the first 10% or 5% of the training set for few-shot training data, this represents a scenario where the model’s forecasting time is far from the training data collection. Thus, we should avoid leaking too much recent data to the model, especially for ICTSP. Because ICTSP can use historical input as in-context training samples, increasing $L_{I}$ too much might leak data that should be invisible to the model, defeating the purpose of the few-shot experiment.
Given the difference between $L_{I}=512$ used by [25, 26] and $L_{I}=1440,L_{b}=512$ for ICTSP, we chose to mask the part of the test set input data not visible to the $L_{I}=512$ setting in [25, 26] with zero-filling for ICTSP. This ensures that no unseen data for [25, 26] is used by ICTSP. As a result, there are no context examples for ICTSP (which reduces it to a Series-wise Transformer) at the very beginning of the test set input data, and the number of context examples gradually increases in the subsequent test set input data points. Thus, the few-shot learning results in Tables 2, 7, and 8 can be regarded as results from a weakened version of ICTSP, yet still achieving SOTA performance.

### v2

@@ 3 Experiments
We conduct comprehensive experiments under full-data, few-shot, and zero-shot settings using widely-used TSF datasets (details in §A.2), including ETTs (Zhou et al., 2021), Traffic, Electricity (ECL), and Weather. We use $K=3$ $\mathtt{TF}$ layers with $d=128$ and 8 heads. We set $L_{I}=1440$, $L_{b}=512$, and $L_{P}\in\{96,192,336,720\}$, performing 4 experiments for each dataset. We use sampling step $m=8$ and the token retrieval method with $q=10\%$, $r=30$ in main experiments. See §A.3.2 for more hyper-parameter and implementation details. We use the same experimental environment as (Zhou et al., 2023; Jin et al., 2024). Extensive efforts are made to ensure fair comparisons across all experiments (detailed in §A.3.3).
Baselines We compare our method with previous SOTA models in the categories of LLM for TSF (Time-LLM (Jin et al., 2024), GPT4TS (Zhou et al., 2023), LLMTime (Gruver et al., 2023)), Temporal-wise Transformers (PatchTST (Nie et al., 2022), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021)), Series-wise Transformer (iTransformer (Liu et al., 2024a)), CNN (TimesNet (Wu et al., 2022)), and simple methods (DLinear (Zeng et al., 2023), Last-value Repeat). Baseline results are sourced from (Zhou et al., 2023; Jin et al., 2024) when applicable and rerun for the missing experiments. We report the average test set MSE for each dataset in the main paper and provide the full raw results in §A.1. The average rank and the number of times each model is ranked as the best are also summarized in the results.
@@
Analysis of TSF Transformer Characteristics From the ICL perspective, the Temporal-wise Transformer is better for learning inter-series relationships, while the Series-wise Transformer better at intra-series relationships. Here, we use a synthesized "Multi" dataset from (Lu et al., 2024) with strong inter-series dependencies, generated by shifting an random walk process. Fig. 4 (a) and (b) show that the Temporal-wise Transformer performs well on this dataset, while the Series-wise Transformer fails to learn the shifting dependencies. ICTSP, with its context tokens featuring a shifting structure, excels in learning these time-interleaved dependencies. Conversely, the Temporal-wise Transformer struggles with real-world datasets with weak dependencies comparing to Series-wise Transformer, such as the ETTm2 datasets in Fig. 4 (c) and (d). ICTSP handles these datasets well, demonstrating significantly better performance. Note that we build these baselines in this analysis with the same structure shown in Fig. 2, using $L_{I}=512$, $L_{P}=192$, $K=3$, $d=128$ for the 3 models; and $L_{b}=256$ for ICTSP. See §A.4.2 for more discussion about the inter-series learning abilities of the TSF Transformers.
@@ 4 Related Works
Extensive research on TSF has been conducted in both traditional statistics (Box et al., 1974; Holt, 2004; Sims, 1980) and deep learning (Hochreiter and Schmidhuber, 1997; Rangapuram et al., 2018; Salinas et al., 2020; Xing et al., 2024). Recent studies have seen a surge in Transformer-based TSF models (Wen et al., 2022; Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Zhang and Yan, 2023). However, a study (Zeng et al., 2023) indicated that these complex models might not outperform linear predictors on real-world datasets, leading to subsequent research (Nie et al., 2022; Liu et al., 2024a; Lu et al., 2024) proposing possible solutions. Additionally, due to the significant achievements of LLMs in various tasks (Brown et al., 2020), several LLM-based TSF models (Gruver et al., 2023; Zhou et al., 2023; Jin et al., 2024) have been introduced. These models also use Transformer architectures but rely on pre-trained parameters and fine-tuning to achieve superior few-shot and zero-shot capabilities (Zhang et al., 2024). While studies (Min et al., 2022; Wei et al., 2023; Xie et al., 2022; Zhang et al., 2023; Garg et al., 2022) have highlighted that these few-shot abilities in LLMs stem from ICL, there is a lack of research on how to efficiently implement ICL in TSF.
@@ A.2 Datasets
ETT Datasets [^2] https://github.com/zhouhaoyi/ETDataset (Zhou et al., 2021): This dataset includes load and oil temperature data from electricity transformers, recorded at 15-minute intervals from July 2016 to July 2018. It comprises four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, representing two transformers (identified as 1 and 2) and two time resolutions (15 minutes and 1 hour). Each subset contains seven features related to oil and load of the transformers.
@@
Multi Dataset (Lu et al., 2024) is generated from a master random walk series. The first series is the master series, while the second to fifth series are created by shifting the master series backward by 96, 192, 336, and 720 steps, respectively. The last three series are combinations of these first five series.
@@ A.3.1 More Details of Model Implementation
The remaining tokens are merged into $r$ tokens by grouped weighted averaging. Let $k=\lfloor qN\rfloor$ be the number of selected tokens, and let p=Nkp=N-k be the number of remaining tokens. The $r$ merged tokens are obtained as follows:
@@ A.3.2 Hyperparameter Settings and Training Details
Model Training Setup The ICTSP model is trained using the Adam optimizer and MSE loss in Pytorch, with a learning rate of 0.0005 each dataset. We test the model every 200 training steps with a early-stopping patience being $30\times 200$ steps. The first 1000 steps are for learning rate warm-up, followed by a linear decay of learning rate. We set the random seed as 2024. Our models are trained on single Nvidia RTX 4090 GPU with a batch size equals to 32 for most of the datasets. We decrease the batch size to 16 and 8 for the larger dataset ECL and Traffic, respectively. In the full-data experiment setting, we split each dataset with 70% training set, 10% validation, set and 20% test set. We fit a standardization scaler on the training set and apply it to the whole dataset. This setting is applied in consistent with previous studies like (Wu et al., 2021; 2022; Nie et al., 2022; Jin et al., 2024).
@@
Randomness of Training We test the randomness of training the ICTSP by alternating the random seed among {2022, 2023, 2024, 2025, 2026} and conducting multiple training runs on the same dataset settings. The error bars of the results with the five random seeds on the Weather and ECL datasets are shown in Figure 8, alongside the results of the two best previous baselines, Time-LLM (Jin et al., 2024) and PatchTST (Nie et al., 2022). The results show that ICTSP can stably surpass previous methods, even with the worst results among the five random seeds, indicating the robustness of ICTSP in performance improvement.
@@ A.3.3 Discussion of Fair Comparison
Our experimental environment aligns with the codes used by previous studies, including (Wu et al., 2021; Zeng et al., 2023; Wu et al., 2022; Nie et al., 2022; Zhou et al., 2023; Jin et al., 2024). Our few-shot and zero-shot learning setups follow (Zhou et al., 2023; Jin et al., 2024).
In the full-data experiment, for the input length $L_{I}$ setting, previous studies like (Nie et al., 2022; Zhou et al., 2023) suggest that $L_{I}$ should be considered a hyper-parameter that needs tuning for each dataset. Therefore, we chose to directly source their results where applicable, to avoid underestimating their performance when rerunning with a fixed $L_{I}$.
In the few-shot experiment, careful consideration of $L_{I}$ is very crucial. Since (Zhou et al., 2023; Jin et al., 2024) use the first 10% or 5% of the training set for few-shot training data, this represents a scenario where the model’s forecasting time is far from the training data collection. Thus, we should avoid leaking too much recent data to the model, especially for ICTSP. Because ICTSP can use historical input as in-context training samples, increasing $L_{I}$ too much might leak data that should be invisible to the model, defeating the purpose of the few-shot experiment.
Given the difference between $L_{I}=512$ used by (Zhou et al., 2023; Jin et al., 2024) and $L_{I}=1440,L_{b}=512$ for ICTSP, we chose to mask the part of the test set input data not visible to the $L_{I}=512$ setting in (Zhou et al., 2023; Jin et al., 2024) with zero-filling for ICTSP. This ensures that no unseen data for (Zhou et al., 2023; Jin et al., 2024) is used by ICTSP. As a result, there are no context examples for ICTSP (which reduces it to a Series-wise Transformer) at the very beginning of the test set input data, and the number of context examples gradually increases in the subsequent test set input data points. Thus, the few-shot learning results in Tables 2, 7, and 8 can be regarded as results from a weakened version of ICTSP, yet still achieving SOTA performance.